# Analytics Modelling over Multiple Datasets using Vector Embeddings

### v3

@@ 3.2 NumTabData2Vec
$NumTabData2Vec\left(D_{i}\right)\rightarrow z,\$ (1)
@@
$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)\$ (2)
@@
$argmax\mathcal{L}_{\theta,\phi}\left(x\right)\$ (4)
@@ 3.3 Dataset Selection
Assign $\text{Assign }\mathbf{z}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$
@@
$SilhouetteScore(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max(a(z_{i}),b(z_{i}))},\$ (5)
where for each vector $z_{i}$ computes the mean intra-cluster distance ($a(z_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(z_{i})$) is the minimum average distance with the other vectors in a different cluster. Silhouette Score ranges from $-1$ to $1$, and the higher value defines the best $s$ number for clusters.
@@ 4 Evaluation
We compared our framework with Apollo [9, 4], which models analytic operators using data content. Two loss functions to measure prediction accuracy are employed: root-mean-square error (RMSE) and mean absolute error (MAE). RMSE is sensitive to outliers, while MAE is not; conversely, RMSE accounts for error direction, which MAE cannot. We further assess efficiency using Speedup and Amortized Speedup metrics, where Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$. Amortized speedup including one-time vectorization per data lake across multiple operators. Three variants with vector sizes 100, 200, and 300 (each with eight transformer layers) were trained for $100$ epochs on four NVIDIA A10101010 GPUs. More experimental evaluation results can be found in the extended version of this work [20].
@@ 4.2 Datasets
Dataset Name # Files # Tuples # Columns
@@ 4.3 Evaluation Results
Dataset Name Method Operator RMSE MAE Speedup Amortized Speedup
Household Power Consumption $300$V Cosine LR $\mathbf{6.61}$ $\mathbf{5.42}$ $0.0017$ $\mathbf{1.99}$
$300$V SR-$0.2$ LR $7.77$ $6.66$ $0.0018$ $1.42$
Apollo-SR $0.1$ LR $2968.01$ $2352.55$ $\mathbf{0.015}$ $0.024$
Apollo-SR $0.2$ LR $2811.49$ $2229.50$ $0.015$ $0.024$
$300$V K-Means MLP Regr. $\mathbf{6.70}$ $\mathbf{3.38}$ $0.9249$ $\mathbf{1.99}$
Apollo-SR $0.1$ MLP Regr. $3322.05$ $2606.99$ $2.38$ $1.74$
Apollo-SR $0.2$ MLP Regr. $3850.01$ $2609.36$ $\mathbf{2.38}$ $1.74$
Stock $300$V Cosine LR $306382.28$ $125335.65$ $0.00085$ $\mathbf{1.91}$
$300$V SR-$0.4$ LR $21861625.91$ $5674215.265$ $0.00087$ $0.33$
Apollo-SR $0.1$ LR $\mathbf{153665.92}$ $\mathbf{118236.48}$ $\mathbf{0.00093}$ $0.00096$
Apollo-SR $0.2$ LR $166844.95$ $133306.68$ $0.00093$ $0.00096$
$300$V Cosine MLP Regr. $\mathbf{140236.47}$ $\mathbf{123571.12}$ $0.63$ $\mathbf{1.91}$
Apollo-SR $0.1$ MLP Regr. $175150.82$ $145123.09$ $\mathbf{0.93}$ $0.96$
Apollo-SR $0.2$ MLP Regr. $174390.81$ $146338.73$ $0.93$ $0.96$
Weather $300$V Cosine SVM SGD $\mathbf{14.13}$ $\mathbf{7.63}$ $1.06$ $\mathbf{22.8}$
Apollo-SR $0.1$ SVM $69.51$ $25.52$ $\mathbf{2.10}$ $1.16$
Apollo-SR $0.2$ SVM $68.70$ $22.81$ $2.10$ $1.16$
$300$V Cosine MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
$300$V SR-$0.4$ MLP $15.95$ $13.31$ $1.02$ $1.77$
Apollo-SR $0.1$ MLP $69.62$ $23.10$ $\mathbf{1.34}$ $1.14$
Apollo-SR $0.2$ MLP $673.56$ $\mathbf{84.70}$ $1.32$ $1.14$
Adult $300$V Cosine SVM SGD $\mathbf{0.36}$ $\mathbf{0.2}$ $0.37$ $\mathbf{2.78}$
Apollo-SR $0.1$ SVM $68.32$ $22.95$ $\mathbf{0.75}$ $0.85$
Apollo-SR $0.2$ SVM $68.88$ $22.88$ $0.74$ $0.85$
$300$V K-Means MLP $\mathbf{0.36}$ $\mathbf{0.19}$ $0.30$ $2.78$
$300$V SR-$0.2$ MLP $6.01$ $6.00$ $0.54$ $\mathbf{3.54}$
Apollo-SR $0.1$ MLP $71.11$ $26.51$ $\mathbf{1.07}$ $1.31$
Apollo-SR $0.2$ MLP $70.16$ $25.74$ $1.05$ $1.31$
@@
Table 2 compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$. Methods $100$V, $200$V, and $300$V denote vector embedding dimensions. The lowest prediction errors align with our pipeline’s similarity search method. Apollo outperforms our framework on the Stock dataset for the LR analytic operator with the smallest amount of SR. However, our framework excels with the MLP regression operator, improving RMSE and MAE by $20\%$ and $17\%$, respectively. The LR operator’s performance gap on the Stock dataset is minor. For other datasets, our framework consistently surpasses Apollo across different SR values. This demonstrates the effectiveness of our similarity search approach, which enhances data quality and reduces $\Phi$ prediction errors by identifying relevant datasets $D_{r}$ from the data lake directory $D$. The Adult dataset also highlights our framework’s advantage with increasing feature dimensions. Although Apollo achieves better raw speedup due to the higher complexity of our framework’s vectorization step, our framework outperforms it in amortized speedup. By excluding the reusable vectorization process, it achieves speed gains of $10\%$ to $60\%$ for most operators. The SR approach, leveraging vector embedding representations, enhances operator prediction compared to Apollo and achieves greater amortized speedup. However, the similarity search method outperforms both Apollo and the SR approach in prediction accuracy and amortized speedup, establishing its clear superiority across most datasets and operator scenarios.
@@ 4.4 NumTabData2Vec Evaluation Results
NumTabData2Vec Vector Size Similarity
$100$ $0.54$
$200$ $0.18$
$300$ $0.16$
@@
# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time $3$ $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec $3$ $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec $3$ $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $3$ $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $20$ $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec $20$ $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec $20$ $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec $20$ $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
@@ 5 Limitations
There exist a number of limitations in our work as we described it. In this section we briefly highlight them. Firstly, our input datasets comprise records of specific size and type (numerical). This currently excludes data with textual and categorical attributes, or tables with varying number of features inside a set of datasets. Secondly, we currently consider single-input and single-output operator modelling. Finally, our proposed NumData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples. This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available for a budget GPU).
@@ 6 Conclusion
In this paper, we presented a novel framework for the modelling of an analytic operator (such as a ML algorithm) when a large number of input data is available and thus no brute-force execution can be performed. We propose a deep learning model, NumData2Vec, which transforms a dataset to a lower $k$-dimensional representation space $z$. Our framework produces vector embeddings for the input datasets using NumData2Vec and performs a similarity search to identify the most relevant subset of datasets for any unseen input. By modelling the analytic operator based on this selected subset, we are able to accurately predict its output on any given input dataset. In practice, we demonstrated that our framework can accurately model various common algorithms and compared favourably against a similar recent framework [9], in both accuracy and speedup. Furthermore, we showed that NumData2Vec can create different vector representations for datasets from different scenarios. We also demonstrated that NumData2Vec can effectively detect when noise is introduced into a dataset.

### v4

@@ 3.2 NumTabData2Vec
$NumTabData2Vec\left(D_{i}\right)\rightarrow z,\\$ (1)
@@
$minD_{KL}\left(q_{\phi}\left(z|x\right)\|p_{\theta}\left(x|z\right)\right)\\$ (2)
@@
$argmax\mathcal{L}_{\theta,\phi}\left(x\right)\\$ (4)
@@ 3.3 Dataset Selection
$\text{Assign }\mathbf{z}_{i}\text{ to cluster }j=\arg\min_{j}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|^{2}$
@@
$SilhouetteScore(z_{i})=\frac{b(z_{i})-a(z_{i})}{\max(a(z_{i}),b(z_{i}))},\\$ (5)
where for each vector $z_{i}$ computes the mean intra-cluster distance ($a(z_{i})$) which is the distance with the other vectors in the same cluster, and the mean nearest-cluster distance ($b(z_{i})$) is the minimum average distance with the other vectors in a different cluster. Silhouette Score ranges from 1-1 to $1$, and the higher value defines the best $s$ number for clusters.
@@ 4 Evaluation
We compared our framework with Apollo [9, 4], which models analytic operators using data content. Two loss functions to measure prediction accuracy are employed: root-mean-square error (RMSE) and mean absolute error (MAE). RMSE is sensitive to outliers, while MAE is not; conversely, RMSE accounts for error direction, which MAE cannot. We further assess efficiency using Speedup and Amortized Speedup metrics, where Speedup is defined as $\frac{T{{}^{(i)}_{op}}}{T{{}^{(i)}_{SimOp}+T_{vec}+T_{sim}+T_{pred}}}$, where $T{{}^{(i)}_{op}}$ is the time to execute operator $i$ on all datasets, $T{{}^{(i)}_{SimOp}}$ is the time to model the operator with datasets from similarity search, $T_{vec}$ is the vector embedding computation time, $T_{sim}$ is the similarity search time, and $T_{pred}$ is the prediction time for $D_{o}$. Amortized speedup including one-time vectorization per data lake across multiple operators. Three variants with vector sizes 100, 200, and 300 (each with eight transformer layers) were trained for $100$ epochs on four NVIDIA A1010 GPUs. More experimental evaluation results can be found in the extended version of this work [20].
@@ 4.2 Datasets
Dataset Name
# Files
# Tuples
# Columns
@@ 4.3 Evaluation Results
Dataset
Name
Method Operator RMSE MAE Speedup Amortized Speedup
Household
Power
Consumption
300300V Cosine
LR $\mathbf{6.61}$ $\mathbf{5.42}$ $0.0017$ $\mathbf{1.99}$
300300V SR-$0.2$
LR $7.77$ $6.66$ $0.0018$ $1.42$
Apollo-SR $0.1$
LR $2968.01$ $2352.55$ $\mathbf{0.015}$ $0.024$
Apollo-SR $0.2$
LR $2811.49$ $2229.50$ $0.015$ $0.024$
300300V K-Means
MLP Regr. $\mathbf{6.70}$ $\mathbf{3.38}$ $0.9249$ $\mathbf{1.99}$
Apollo-SR $0.1$
MLP Regr. $3322.05$ $2606.99$ $2.38$ $1.74$
Apollo-SR $0.2$
MLP Regr. $3850.01$ $2609.36$ $\mathbf{2.38}$ $1.74$
Stock
300300V Cosine
LR $306382.28$ $125335.65$ $0.00085$ $\mathbf{1.91}$
300300V SR-$0.4$
LR $21861625.91$ $5674215.265$ $0.00087$ $0.33$
Apollo-SR $0.1$
LR $\mathbf{153665.92}$ $\mathbf{118236.48}$ $\mathbf{0.00093}$ $0.00096$
Apollo-SR $0.2$
LR $166844.95$ $133306.68$ $0.00093$ $0.00096$
300300V Cosine
MLP Regr. $\mathbf{140236.47}$ $\mathbf{123571.12}$ $0.63$ $\mathbf{1.91}$
Apollo-SR $0.1$
MLP Regr. $175150.82$ $145123.09$ $\mathbf{0.93}$ $0.96$
Apollo-SR $0.2$
MLP Regr. $174390.81$ $146338.73$ $0.93$ $0.96$
Weather
300300V Cosine
SVM SGD
$\mathbf{14.13}$ $\mathbf{7.63}$ $1.06$ $\mathbf{22.8}$
Apollo-SR $0.1$
SVM $69.51$ $25.52$ $\mathbf{2.10}$ $1.16$
Apollo-SR $0.2$
SVM $68.70$ $22.81$ $2.10$ $1.16$
300300V Cosine
MLP $\mathbf{14.29}$ $\mathbf{4.03}$ $1.03$ $\mathbf{22.8}$
300300V SR-$0.4$
MLP $15.95$ $13.31$ $1.02$ $1.77$
Apollo-SR $0.1$
MLP $69.62$ $23.10$ $\mathbf{1.34}$ $1.14$
Apollo-SR $0.2$
MLP $673.56$ $\mathbf{84.70}$ $1.32$ $1.14$
Adult
300300V Cosine
SVM SGD
$\mathbf{0.36}$ $\mathbf{0.2}$ $0.37$ $\mathbf{2.78}$
Apollo-SR $0.1$
SVM $68.32$ $22.95$ $\mathbf{0.75}$ $0.85$
Apollo-SR $0.2$
SVM $68.88$ $22.88$ $0.74$ $0.85$
300300V K-Means
MLP $\mathbf{0.36}$ $\mathbf{0.19}$ $0.30$ $2.78$
300300V SR-$0.2$
MLP $6.01$ $6.00$ $0.54$ $\mathbf{3.54}$
Apollo-SR $0.1$
MLP $71.11$ $26.51$ $\mathbf{1.07}$ $1.31$
Apollo-SR $0.2$
MLP $70.16$ $25.74$ $1.05$ $1.31$
@@
Table 2 compares model operators, loss functions, and speedup metrics for our framework and Apollo at SR values of $0.1$ and $0.2$. Methods 100100V, 200200V, and 300300V denote vector embedding dimensions. The lowest prediction errors align with our pipeline’s similarity search method. Apollo outperforms our framework on the Stock dataset for the LR analytic operator with the smallest amount of SR. However, our framework excels with the MLP regression operator, improving RMSE and MAE by $20\%$ and $17\%$, respectively. The LR operator’s performance gap on the Stock dataset is minor. For other datasets, our framework consistently surpasses Apollo across different SR values. This demonstrates the effectiveness of our similarity search approach, which enhances data quality and reduces $\Phi$ prediction errors by identifying relevant datasets $D_{r}$ from the data lake directory $D$. The Adult dataset also highlights our framework’s advantage with increasing feature dimensions. Although Apollo achieves better raw speedup due to the higher complexity of our framework’s vectorization step, our framework outperforms it in amortized speedup. By excluding the reusable vectorization process, it achieves speed gains of $10\%$ to $60\%$ for most operators. The SR approach, leveraging vector embedding representations, enhances operator prediction compared to Apollo and achieves greater amortized speedup. However, the similarity search method outperforms both Apollo and the SR approach in prediction accuracy and amortized speedup, establishing its clear superiority across most datasets and operator scenarios.
@@ 4.4 NumTabData2Vec Evaluation Results
NumTabData2Vec
Vector Size
Similarity
$100$
$0.54$
$200$
$0.18$
$300$
$0.16$
@@
# of columns # of rows $50$ Vectors Execution time $100$ Vectors Execution time $200$ Vectors Execution time 33 $100$ $0.0004$ sec $0.00042$ sec $0.00051$ sec 33 $500$ $0.0004$ sec $0.00041$ sec $0.00049$ sec 33 $1000$ $0.0004$ sec $0.00041$ sec $0.00049$ sec 33 $1500$ $0.0004$ sec $0.00041$ sec $0.00055$ sec 33 $1800$ $0.0004$ sec $0.00041$ sec $0.00055$ sec $10$ $100$ $0.0004$ sec $0.0004$ sec $0.00057$ sec $10$ $500$ $0.00039$ sec $0.0004$ sec $0.00051$ sec $10$ $1000$ $0.00041$ sec $0.00042$ sec $0.00052$ sec $10$ $1500$ $0.00041$ sec $0.00042$ sec $0.00055$ sec $10$ $1800$ $0.00041$ sec $0.00042$ sec $0.00052$ sec 2020 $100$ $0.0004$ sec $0.00042$ sec $0.0005$ sec 2020 $500$ $0.0004$ sec $0.00042$ sec $0.0005$ sec 2020 $1000$ $0.00042$ sec $0.00043$ sec $0.00052$ sec 2020 $1500$ $0.00043$ sec $0.00044$ sec $0.00054$ sec 2020 $1800$ $0.00044$ sec $0.00044$ sec $0.00054$ sec
@@ 5 Limitations
There exist a number of limitations in our work as we described it. In this section we briefly highlight them. Firstly, our input datasets comprise records of specific size and type (numerical). This currently excludes data with textual and categorical attributes, or tables with varying number of features inside a set of datasets. Secondly, we currently consider single-input and single-output operator modelling. Finally, our proposed NumTabData2Vec model for data vectorization has a performance limitation, as it cannot deal with datasets bigger than about $3000$ tuples. This is mostly a hardware limitation of off-the-shelf GPUs (with at most 24GB of memory available for a budget GPU).
@@ 6 Conclusion
In this paper, we presented a novel framework for the modelling of an analytic operator (such as a ML algorithm) when a large number of input data is available and thus no brute-force execution can be performed. We propose a deep learning model, NumTabData2Vec, which transforms a dataset to a lower $k$-dimensional representation space $z$. Our framework produces vector embeddings for the input datasets using NumTabData2Vec and performs a similarity search to identify the most relevant subset of datasets for any unseen input. By modelling the analytic operator based on this selected subset, we are able to accurately predict its output on any given input dataset. In practice, we demonstrated that our framework can accurately model various common algorithms and compared favourably against a similar recent framework [9], in both accuracy and speedup. Furthermore, we showed that NumTabData2Vec can create different vector representations for datasets from different scenarios. We also demonstrated that NumTabData2Vec can effectively detect when noise is introduced into a dataset.